An alloy of tin is 16% tin. How much tin must be added to make 820 lbs of alloy that is 18% tin? Answer to the nearest tenth.
Answer by checkley71(8403)

.16X+(820-X)=.18*820
.16X+820-X=147.6
-.84X=147.6-820
-.84=-672.4
X=-672.4/-.84
X=800.5 ANSWER FOR THE AMOUNT OF 18% TIN.
820-800=20 LBS. OF PURE TIN IS USED.
PROOF.
.16*800.5+19.5=.18*820
128.08+19.5=147.6
147.6~147.6